# A Cascaded Random Access Quantum Memory

### v1

@@ Appendix F Fast Flux Delivery for High-Q 3D Cavities
(1) Ultra-wide ($>3$ GHz) and deep stopband with a flexible flux line geometry.
@@
(4) Free from external flux coils while maintaining high cavity coherence ($>1$ ms) and low cavity thermal populations ($<0.05\%$).
Our coupler chip design is shown in Fig. 10(a). It is composed of 4 parts (shown in Fig. 10(b), (c), (d), (e)): The SQUID coupler, the shorting lines, the edge-coupled microstrip filter, and the double-Y balun.
The SQUID coupler is a flux-tunable transmon with two pads that are capacitively coupled to both buffer and storage modes. The flux loop is designed to deliver a pure RF flux modulation to the SQUID. However, the flux loop also has a stray capacitive coupling to the coupler. One solution is to maximize the flux line’s inductive coupling to the SQUID loop [45, 46]. In these references involving 2D designs, the flux line produced by such optimization is effective for all frequency ranges and helps mediate fast and clean two-mode squeezing sidebands at $>7\,\mathrm{GHz}$. In this work, we can alternatively optimize the SQUID coupler geometries for low RF modulation frequencies $<2\,\mathrm{GHz}$ [27, 47]. We optimized the location of the flux loop using ANSYS HFSS simulations. In particular, we quantified the flux-to-charge drive ratio with the following equations: We define the two SQUID junctions as $J_{1}$ and $J_{2}$. When RF signals at frequency $\omega_{rf}$ are applied through the flux line, the induced currents along $J_{1}$ and $J_{2}$ are denoted $I_{1}(\omega_{rf})$ and $I_{2}(\omega_{rf})$, respectively [27]. For a pure flux drive:
@@
Shorting lines are for ESD protection of fab on sapphire chips. They are removed later during packaging and are discussed in the Supplementary Section. J.
The edge-coupled microstrip filter [48] provides a wide stopband to protect all coherent elements in our RAQM. As shown in Fig. 11(c), the simulated response (blue curve) demonstrates this stopband. Similar to a Step-Impedance Low-Pass Filter (SIPF), the edge-coupled microstrip filter consists of alternating high- and low-impedance coplanar stripline (CPS) sections of equal electrical length. The stopband’s width, depth, and center frequency are determined independently by the impedance ratio, the number of sections, and the electrical length. The analytic expression [49] can be used to quickly calculate the ABCD matrix of the filter and approximate the stopband shape.
@@
The flux filter works effectively when the input impedance is constant across all $\omega_{rf}$. When coaxial cables transmit flux signals through the flux line, the double-Y balun helps suppress filter ripples caused by impedance mismatch on the package side. Fig. 11(a) and (b) show HFSS simulations for two cases: a perfect on-chip $50\,\Omega$ excitation port, and the entire package. The resulting RF modulation responses, depicted in Fig. 11(c) by the black and blue curves, demonstrate that the double-Y balun effectively aligns the two curves at low frequencies. More significant ripples appear in the blue curve at frequencies ($>7$ GHz) higher than the balun range. Reducing the size of the double-Y balun can rapidly increase the cutoff frequency. In our design, the balun size ($\sim 0.5$ mm) is constrained by the need for additional wirebonds. For two-mode squeezing interactions with a frequency around $10$ GHz, directly fabricating airbridges on the balun and optimizing the design of CPS and CPW open ports will help achieve the desired frequency range.
@@ G.1 State-Dependent Access Error
Figure 12: RB for state-dependent access error. RB is performed in the target mode $S_{i}$ to benchmark the $BS_{i}$ swap fidelity while the spectator mode $S_{j}$ is occupied. The decay curve for each RB (target $S_{i}$, spectator $S_{j}$) is additionally averaged over $6$ cases, one for each of the possible spectator occupation: $\{\left|0\right\rangle,\left|1\right\rangle,\left|+x\right\rangle,\left|-x\right\rangle,\left|+y\right\rangle,\left|-y\right\rangle\}$. (a) Semiclassical fit for the $(S_{i},S_{j})=(S_{4},S_{6})$ RB, reporting the additional swap error $\varepsilon_{i}^{j}=0.938\pm 0.060\%$. (b) Benchmarking $B_{1}S_{i}=B_{1}S_{4}$ swap fidelity in the presence of random population in $S_{j}$. The reference data refers to no population in the spectator mode.
The state-dependent access error occurs when fetching the target mode via the $BS_{i}$ swap. When the other storage modes $S_{j},(j\neq i)$ are occupied, the buffer-storage cross-Kerr interactions $\chi_{BS_{j}}$ can change the $BS_{i}$ swap frequency randomly. Those randomly applied off-resonant swaps have lower fidelity and finally dephase the state in $S_{i}$ after rounds of read/write. To benchmark this error, we perform randomized benchmarking (RB) introduced in the main text Section III between $B_{1}$ and $S_{i}$ in the presence of populations in distinct spectator mode $S_{j}$. During the RAM RB experiments, the RAQM is populated with different Clifford states in the $\{\left|0\right\rangle,\left|1\right\rangle\}$ subspace. We thus perform $6$ rounds of RB to check the $BS_{i}$ swap fidelity with the $6$ distinct Clifford states in $S_{j}$. The final RB decay curve is average over the $6$ curves. We evaluated fidelities for $15$ different gate depths in the range $[0,500]$. For each depth, we chose $20$ different sequences for each RB gate depth, with each sequence measured $1000-5000$ times depending on gate depth (more repetitions for larger gate depths).
To extract the state-dependent access error, we perform a semi-classical Markovian analysis based on the measured RB decay curve. This analysis captures the effect of decaying spectator populations, which causes the state-dependent access error to be lower for large gate depths. As shown in Figure 12 (b), after $300$ gates (${}450\mu$s), decay curves for all target-spectator $S_{i}S_{j}$ pairs follow the same trend as the reference case. We will now outline the analysis procedure.
@@
To extract $\varepsilon_{i}^{j}$, we use the above equation to fit the previously obtained decay curve to a depth up to $50$ gates ($\sim 50-75\mu\mathrm{s}$). We set $\zeta=1/2$ as the decay curve averages over the spectator mode’s population over the $6$ cardinal states of its Bloch sphere. An example of this fitting is shown in Fig. 12 (a). The state-dependent access error for a swap gate is $\varepsilon_{ai}^{j}=(1-F_{0}^{i}\varepsilon_{i}^{j})^{3/2}$, which is plotted in Fig 4(c). For most target-spectator mode pairs, the state-dependent access error is below $<0.3\%$, while only a few pairs have an error rate as high as $1.396\%$.
@@ G.3 Spectator Access Dephasing
As shown in Fig. 4 (d), the spectator access dephasing rate is below $3\,\mathrm{kHz}$ for most target-spectator pairs, while some have as high as $18.75\,\mathrm{kHz}$. As RAQM size increases, each mode spends more time idling per memory flashing round. As a result, the modes become more sensitive to spectator access dephasing as we scale the size of RAQM.
@@ Appendix H Single Mode Control
After the swap, the buffer state $x\left|0\right\rangle+y\left|1\right\rangle$ is mapped to the transmon state $x\left|g\right\rangle+ye^{1j\phi_{d4}}\left|1\right\rangle$. Performing a single transmon gate requires a virtual phase correction to reverse the deterministic phase $\phi_{d4}$ coming from the sideband stark-shift. The phase cancellation procedure is similar to that described in Supplementary Section E. After performing the single transmon gate, the state is swapped back to the buffer. Fig. 13 shows the single transmon gate fidelity ($99.958\%\pm 0.003\%$) measured through RB.
The single buffer gate length is $1.65\,\mu\mathrm{s}$ with an average Clifford gate fidelity of $98.64\%$. This includes two $0.615\,\mu\mathrm{s}$ buffer-transmon swaps, two $140\,\mu\mathrm{s}$ transmon $\left|f\right\rangle\leftrightarrow\left|e\right\rangle$ $\pi$ pulses, and a $140\,\mu\mathrm{s}$ transmon $\left|e\right\rangle\leftrightarrow\left|g\right\rangle$ pulse. Compared to the small transmon-buffer $\chi$-shift ($1/\chi_{qb}\sim 3.5\,\mu\mathrm{s}$), this is a cavity gate scheme beyond the $\chi$-limit. Our experimental gate speed is limited by the room-temperature amplifier power on the drive lines.
@@
Appendix I Error analysis
I.1 Single-mode error budget
@@
Fig. 14 shows the error budget for the swap fidelity of all $B_{1}S_{i}$ pairs. The raw and post-select fidelities are all experimentally measured through randomized benchmarking. The blue and green parts represent the $T_{1}$ decay error from the $B_{1}$ and $S_{i}$ during the swap. We use the $B_{1}$ and $S_{i}$’s coherence during idle time (Table. 1) to calculate the decay error contribution during beamsplitter interactions. The remaining mismatch in fidelities could come from modes’ $T_{1}$ fluctuations. The post-selected fidelity may be limited by coupler heating and cavity dephasing.
I.2 Full RAQM error budget
@@
Here, we discuss the error budget for random read infidelity. Fig. 15 shows the pulse sequence for each $S_{i}$’s access gate. Here, we label the gate lengths for a $B_{1}S_{i}$ swap, a $QB_{1}$ $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ swap, a transmon $\left|e\right\rangle\leftrightarrow\left|f\right\rangle$ $\pi$ pulse, and a random transmon gate $G$ on the $\{\left|g\right\rangle,\left|e\right\rangle\}$ subspace as $2t_{BS_{i}}$, $t_{f0g1}$, $t_{ef}$, and $t_{ge}$. While $t_{BS_{i}}$ are dependent on $S_{i}$, the other gate lengths are fixed in our experiments: $t_{f0g1}=0.606\,\mu\mathrm{s}$ and $t_{ef}=t_{ge}=0.14\,\mu\mathrm{s}$. Each $S_{i}$’s access gate is divided into read, transmon processing, and write periods with length $t_{\text{read}}$, $t_{\text{write}}$, and $t_{\text{proc}}$. Information is transferred from $S_{i}$ to $B_{1}$ during the read period, while the write period reverses this process. Under this definition, we have:
@@
The coefficient accounts for the number of storage modes ($7$) in the multiplexed RAQM operation to avoid double counting (the same for the rest of the error budget discussions). Table 6 shows the experimentally measured state-dependent access error.
@@
Details about the spectator-access dephasing rate are shown in Table 7. For each pair’s RB, we use $30$ different sequences for each RB gate depth to get statistics.
@@
State-dependent access error $0.164\%$ $0.245\%$ $0.058\%$ $0.255\%$ $0.056\%$ $0.167\%$ $0.127\%$
Many-body dephasing $0.102\%$ $0.046\%$ $0.030\%$ $0.026\%$ $0.030\%$ $0.028\%$ $0.042\%$
@@
$\Gamma_{1}^{i}$ and $\Gamma_{\phi}^{i}$ represent the decay and pure dephasing rate for $S_{i}$. Table. 8 lists the error budget in detail. The average random read fidelities of different RAQM sizes are shown in Table. 9.
@@
I.3 Master equation simulation
In this section, we perform a master-equation simulation to study the impact of decoherence on the fidelity of $BS_{i}$ swap gates using QuTiP  [53]. We focused on $BS_{2}$ gate, which had the lowest infidelity as shown in Fig. 2(e). We considered the following lab-frame Hamiltonian:
@@ I.3 Master equation simulation
[…]
@@
Dephasing $[\Gamma_{\phi,b},\Gamma_{\phi,s}]$ $0.1095\%\pm 0.0019\%$ $0.0481\%\pm 0.0005\%$
Decay $[\Gamma_{1,b},\Gamma_{1,s},\Gamma_{1,c}]$ $0.4007\%\pm 0.0043\%$ $0.0039\%\pm 0.0001\%$
Decay and Dephasing $[\Gamma_{1,b},\Gamma_{1,s},\Gamma_{1,c},\Gamma_{\phi,b},\Gamma_{\phi,s}]$ $0.4408\%\pm 0.0052\%$ $0.0479\%\pm 0.0006\%$
Experiment $0.35789\%\pm 0.00002\%$ $0.036890\%\pm 0.01520\%$
@@
Appendix J Electrostatic Discharge Protection
@@
Appendix K Device Fabrication and Measurement Setup

### v2

@@ Appendix F Fast Flux Delivery for High-Q 3D Cavities
(1) Ultra-wide (>3>3 GHz) and deep stopband with a flexible flux line geometry.
@@
(4) Free from external flux coils while maintaining high cavity coherence (>1>1 ms) and low cavity thermal populations ($0.05\%$).
Our coupler chip design is shown in Supplementary Figure 10(a). It is composed of 4 parts (shown in Supplementary Figure 10(b), (c), (d), (e)): The SQUID coupler, the shorting lines, the edge-coupled microstrip filter, and the double-Y balun.
The SQUID coupler is a flux-tunable transmon with two pads that are capacitively coupled to both buffer and storage modes. The flux loop is designed to deliver a pure RF flux modulation to the SQUID. However, the flux loop also has a stray capacitive coupling to the coupler. One solution is to maximize the flux line’s inductive coupling to the SQUID loop [37, 36]. In these references involving 2D designs, the flux line produced by such optimization is effective for all frequency ranges and helps mediate fast and clean two-mode squeezing sidebands at $7\,\mathrm{GHz}$. In this work, we can alternatively optimize the SQUID coupler geometries for low RF modulation frequencies $2\,\mathrm{GHz}$ [39, 38]. We optimized the location of the flux loop using ANSYS HFSS simulations. In particular, we quantified the flux-to-charge drive ratio with the following equations: We define the two SQUID junctions as $J_{1}$ and $J_{2}$. When RF signals at frequency $\omega_{rf}$ are applied through the flux line, the induced currents along $J_{1}$ and $J_{2}$ are denoted $I_{1}(\omega_{rf})$ and $I_{2}(\omega_{rf})$, respectively [39]. For a pure flux drive:
@@
Shorting lines are for ESD protection of fab on sapphire chips. They are removed later during packaging and are discussed in the Supplementary Section. K.
The edge-coupled microstrip filter [67] provides a wide stopband to protect all coherent elements in our RAQM. As shown in Supplementary Figure 11(c), the simulated response (blue curve) demonstrates this stopband. Similar to a Step-Impedance Low-Pass Filter (SIPF), the edge-coupled microstrip filter consists of alternating high- and low-impedance coplanar stripline (CPS) sections of equal electrical length. The stopband’s width, depth, and center frequency are determined independently by the impedance ratio, the number of sections, and the electrical length. The analytic expression [29] can be used to quickly calculate the ABCD matrix of the filter and approximate the stopband shape.
@@
The flux filter works effectively when the input impedance is constant across all $\omega_{rf}$. When coaxial cables transmit flux signals through the flux line, the double-Y balun helps suppress filter ripples caused by impedance mismatch on the package side. Supplementary Figure 11(a) and (b) show HFSS simulations for two cases: a perfect on-chip $50\,\Omega$ excitation port, and the entire package. The resulting RF modulation responses, depicted in Supplementary Figure 11(c) by the black and blue curves, demonstrate that the double-Y balun effectively aligns the two curves at low frequencies. More significant ripples appear in the blue curve at frequencies (>7>7 GHz) higher than the balun range. Reducing the size of the double-Y balun can rapidly increase the cutoff frequency. In our design, the balun size ($\sim 0.5$ mm) is constrained by the need for additional wirebonds. For two-mode squeezing interactions with a frequency around $10$ GHz, directly fabricating airbridges on the balun and optimizing the design of CPS and CPW open ports will help achieve the desired frequency range.
@@ G.1 State-Dependent Access Error
Figure 12: RB for state-dependent access error. RB is performed in the target mode $S_{i}$ to benchmark the $BS_{i}$ swap fidelity while the spectator mode $S_{j}$ is occupied. The decay curve for each RB (target $S_{i}$, spectator $S_{j}$) is additionally averaged over 66 cases, one for each of the possible spectator occupation: $\{\left|0\right\rangle,\left|1\right\rangle,\left|+x\right\rangle,\left|-x\right\rangle,\left|+y\right\rangle,\left|-y\right\rangle\}$. (a) Semiclassical fit for the $(S_{i},S_{j})=(S_{4},S_{6})$ RB, reporting the additional swap error $\varepsilon_{i}^{j}=0.938\pm 0.060\%$. (b) Benchmarking $B_{1}S_{i}=B_{1}S_{4}$ swap fidelity in the presence of random population in $S_{j}$. The reference data refers to no population in the spectator mode.
The state-dependent access error occurs when fetching the target mode via the $BS_{i}$ swap. When the other storage modes $S_{j},(j\neq i)$ are occupied, the buffer-storage cross-Kerr interactions $\chi_{BS_{j}}$ can change the $BS_{i}$ swap frequency randomly. Those randomly applied off-resonant swaps have lower fidelity and finally dephase the state in $S_{i}$ after rounds of read/write. To benchmark this error, we perform randomized benchmarking (RB) introduced in the main text Section III between $B_{1}$ and $S_{i}$ in the presence of populations in distinct spectator mode $S_{j}$. During the RAM RB experiments, the RAQM is populated with different Clifford states in the $\{\left|0\right\rangle,\left|1\right\rangle\}$ subspace. We thus perform $6$ rounds of RB to check the $BS_{i}$ swap fidelity with the $6$ distinct Clifford states in $S_{j}$. The final RB decay curve is average over the $6$ curves. We evaluated fidelities for $15$ different gate depths in the range $[0,500]$. For each depth, we chose $20$ different sequences for each RB gate depth, with each sequence measured 100050001000-5000 times depending on gate depth (more repetitions for larger gate depths).
[…]
@@
To extract $\varepsilon_{i}^{j}$, we use the above equation to fit the previously obtained decay curve to a depth up to $50$ gates ($\sim 50-75\mu\mathrm{s}$). We set $\zeta=1/2$ as the decay curve averages over the spectator mode’s population over the $6$ cardinal states of its Bloch sphere. An example of this fitting is shown in Supplementary Figure 12 (a). The state-dependent access error for a swap gate is $\varepsilon_{ai}^{j}=(1-F_{0}^{i}\varepsilon_{i}^{j})^{3/2}$, which is plotted in the main text Fig. 4 (c). For most target-spectator mode pairs, the state-dependent access error is below $0.3\%$, while only a few pairs have an error rate as high as $1.396\%$.
@@ G.3 Spectator Access Dephasing
As shown in the main text Fig. 4 (d), the spectator access dephasing rate is below $3\,\mathrm{kHz}$ for most target-spectator pairs, while some have as high as $18.75\,\mathrm{kHz}$. As RAQM size increases, each mode spends more time idling per memory flashing round. As a result, the modes become more sensitive to spectator access dephasing as we scale the size of RAQM.
@@ Appendix H Single Mode Control
After the swap, the buffer state $x\left|0\right\rangle+y\left|1\right\rangle$ is mapped to the transmon state $x\left|g\right\rangle+ye^{1j\phi_{d4}}\left|1\right\rangle$. Performing a single transmon gate requires a virtual phase correction to reverse the deterministic phase $\phi_{d4}$ coming from the sideband stark-shift. The phase cancellation procedure is similar to that described in Supplementary Section E. After performing the single transmon gate, the state is swapped back to the buffer. Supplementary Figure 13 shows the single transmon gate fidelity ($99.958\%\pm 0.003\%$) measured through RB.
The single buffer gate length is $1.65\,\mu\mathrm{s}$ with an average Clifford gate fidelity of $98.64\%$. This includes two $0.615\,\mu\mathrm{s}$ buffer-transmon swaps, two $140\,\mathrm{ns}$ transmon $\left|f\right\rangle\leftrightarrow\left|e\right\rangle$ $\pi$ pulses, and a $140\,\mathrm{ns}$ transmon $\left|e\right\rangle\leftrightarrow\left|g\right\rangle$ pulse. Compared to the small transmon-buffer $\chi$-shift ($1/\chi_{qb}\sim 3.5\,\mu\mathrm{s}$), this is a cavity gate scheme beyond the $\chi$-limit. Our experimental gate speed is limited by the room-temperature amplifier power on the drive lines.
@@
Appendix I Random Read Fidelity
The Random read fidelity $\bar{F}_{BS_{i}}$ is measured through multiplexed RB across all memory modes. For each mode $S_{i}$, the depolarization parameter $P_{S_{i}}$ is extracted by fitting the process fidelity after different RAQM cycles. For a given mode $S_{i}$, a single cycle in a size-7 memory consists of 14 gates: two gates (read and write) are performed on $S_{i}$, while the other 12 gates are performed on the other storage modes. The latter are effectively identity gates on $S_{i}$. However, due to decoherence and modes’ cross-Kerrs, these identity gates have nonzero infidelities. $S_{i}$’s random read fidelity $\bar{F}_{BS_{i}}$ is
Here, we isolate the buffer mode’s control fidelity, with its depolarization parameter defined as $P_{B}$. $P_{B}$ is measured independently using single-mode RB on the buffer, involving only $G_{i}^{n}$. The factor of $2/3$ arises because three levels are used in the RAQM controlling transmon [30], and the transmon $\pi_{ef}$ rotation spreads error population from $\left|e\right\rangle$ to $\left|f\right\rangle$. This factor thus provides a conservative estimate of the fidelity.
Appendix J Error analysis
J.1 Single-mode error budget
@@ Appendix I Random Read Fidelity
Supplementary Figure 14 shows the error budget for the swap fidelity of all $B_{1}S_{i}$ pairs. The raw and post-select fidelities are all experimentally measured through randomized benchmarking. The blue and green parts represent the $T_{1}$ decay error from the $B_{1}$ and $S_{i}$ during the swap. We use the $B_{1}$ and $S_{i}$’s coherence during idle time (Supplementary Table 1) to calculate the decay error contribution during beamsplitter interactions. The remaining mismatch in fidelities could come from modes’ $T_{1}$ fluctuations. The post-selected fidelity may be limited by coupler heating and cavity dephasing.
J.2 Full RAQM error budget
@@
Here, we discuss the error budget for random read infidelity. Supplementary Figure 15 shows the pulse sequence for each $S_{i}$’s access gate. Here, we label the gate lengths for a $B_{1}S_{i}$ swap, a $QB_{1}$ $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ swap, a transmon $\left|e\right\rangle\leftrightarrow\left|f\right\rangle$ $\pi$ pulse, and a random transmon gate $G$ on the $\{\left|g\right\rangle,\left|e\right\rangle\}$ subspace as $2t_{BS_{i}}$, $t_{f0g1}$, $t_{ef}$, and $t_{ge}$. While $t_{BS_{i}}$ are dependent on $S_{i}$, the other gate lengths are fixed in our experiments: $t_{f0g1}=0.606\,\mu\mathrm{s}$ and $t_{ef}=t_{ge}=0.14\,\mu\mathrm{s}$. Each $S_{i}$’s access gate is divided into read, transmon processing, and write periods with length $t_{\text{read}}$, $t_{\text{write}}$, and $t_{\text{proc}}$. Information is transferred from $S_{i}$ to $B_{1}$ during the read period, while the write period reverses this process. Under this definition, we have:
@@
The coefficient accounts for the number of storage modes ($7$) in the multiplexed RAQM operation to avoid double counting (the same for the rest of the error budget discussions). Supplementary Table 6 shows the experimentally measured state-dependent access error.
@@
Details about the spectator-access dephasing rate are shown in Supplementary Table 7. For each pair’s RB, we use $30$ different sequences for each RB gate depth to get statistics.
@@
State-dependent access error $0.164\%$ $245\% 0.0$$058\%$ $255\% 0.0$$056\%$ $0.167\%$ $0.127\%$
Many-body dephasing $102\% 0.0$$46\% 0.0$$30\% 0.0$$26\% 0.0$$30\% 0.0$$28\% 0.0$$042\%$
@@
$\Gamma_{1}^{i}$ and $\Gamma_{\phi}^{i}$ represent the decay and pure dephasing rate for $S_{i}$. Supplementary Table 8 lists the error budget in detail. The average random read fidelities of different RAQM sizes are shown in Supplementary Table 9.
@@
J.3 Master equation simulation
In this section, we perform a master-equation simulation to study the impact of decoherence on the fidelity of $BS_{i}$ swap gates using QuTiP  [27]. We focused on $BS_{2}$ gate, which had the lowest infidelity as shown in the main text Fig. 2 (e). We considered the following lab-frame Hamiltonian:
@@
To calibrate the beamsplitter gate $U_{\text{b.s}}(0)=X/2$, we first selected the drive amplitude and frequency which maximized $\mathcal{F}=|\langle 0_{L}|1_{L}\rangle|^{2}$ after a $X/2+X/2$ pulse sequence. Each $X/2$ is a flat top pulse with flat length $0.532\mathrm{\,\mu s}$ and a gaussian ramp of duration $15\text{ ns}$ and $\sigma=5\text{ ns}$, which are the same parameters as that used in the experiment. Since the drive amplitude and frequency have a nonlinear relationship, we swept over both parameters to arrive at the optimal point ($\epsilon=0.03505$, $\omega=514.1731\,\mathrm{MHz}$) which yielded $1-\mathcal{F}<10^{-5}$. Moreover, we added a global phase to convert the $\sqrt{\text{iSWAP}}$ into $\sqrt{\text{SWAP}}$ gate.
@@
Dephasing $\Gamma_{\phi,b},\Gamma_{\phi,s}]$ $1095\%\pm 0.0019\% 0.0$$0481\%\pm 0.0005\%$
Decay $\Gamma_{1,b},\Gamma_{1,s},\Gamma_{1,c}]$ $4007\%\pm 0.0043\% 0.0$$0039\%\pm 0.0001\%$
Decay and Dephasing $\Gamma_{1,b},\Gamma_{1,s},\Gamma_{1,c},\Gamma_{\phi,b},\Gamma_{\phi,s}]$ $4408\%\pm 0.0052\% 0.0$$0479\%\pm 0.0006\%$
Experiment $35789\%\pm 0.00002\% 0.0$$036890\%\pm 0.01520\%$
@@
Appendix K Electrostatic Discharge Protection
@@
Appendix L RAQM Advantage
We define the RAQM advantage as the reduction of control lines in a quantum computer without compromising the number of logical qubits and the logical error rate. Ideally, when all memory modes have infinitely long coherence and the memory read/write error rate is less than the error correction threshold, we can control an arbitrary number of logical qubits with a single set of control lines. Increasing the RAQM size here incurs no extra cost, as no errors accumulate while information stays in the RAQM. Hence RAQM advantage exists intuitively.
We now analytically demonstrate this advantage when a RAQM is attached to each data qubit of a given logical encoding, as shown in Fig. 1a in the main text. We define $[n,k,d]$ as the number of physical qubits $n$ (including ancilla qubits for stabilizer measurement), number of logical qubits $k$, and code distance $d$. We further assume that the quantum computer requires $k_{0}$ logical qubits with a logical error rate at most $\varepsilon_{\text{FT}}$. Additionally, we assume that each physical qubit requires $C_{q}$ control lines to perform error correction without RAQM, and $C_{q}+1$ control lines to perform error correction with RAQM.
Without RAQM attachments, we assume that each logical module (set of data and ancilla qubits in the processor layer) employs the error correction code $n_{1},k_{1},d_{1}]$. Here we need $k_{0}/k_{1}$ copies of the error correction code. The total number of control lines is then
$\left(\frac{k_{0}}{k_{1}}\right)\times n_{1}C_{q}.$ (35)
With a RAQM of size $N$ attached to each data qubit, we assume that each logical module employs a bigger error correction code $n_{2},k_{2},d_{2}]$ to achieve the target logical error rate $\varepsilon_{\text{FT}}$. Since incorporating a memory will introduce additional errors, the code distance in this case is no smaller than in the memory-less case: $d_{2}\geq d_{1}$. Using the RAQM, we only need $k_{0}/(k_{2}N)$ copies of the error correction code. Thus, the total number of lines to control $k_{0}$ logical qubits becomes
$\left(\frac{k_{0}}{Nk_{2}}\right)\times n_{2}(C_{q}+1).$ (36)
The RAQM advantage exists if Eq. 36 is smaller than Eq. 35 and the code $n_{2},k_{2},d_{2}]$ enabling fault tolerance with RAQM exists.
We consider two classes of error-correcting codes combined with RAQM: (1) codes with a constant encoding ratio, and (2) the surface code. The first case represents the most efficient form of logical encoding (favorable to RAQM), while the second case is a currently existing code (unfavorable to RAQM). If the RAQM advantage exists in both cases, then any code with an encoding ratio lying between them should also exhibit a regime of RAQM advantage.
L.1 Case 1: Constant Encoding Ratio
In this idealized case, even though incorporating memory requires a larger code distance, the encoding ratio remains the same as in the memory-less case: $k_{1}/n_{1}=k_{2}/n_{2}$. Comparing Eq. 35 and Eq. 36, if a RAQM with $N>2$ is introduced and the combined processor and memory read/write errors per logical cycle remain below threshold, a RAQM advantage exists. For example, good quantum low-density parity-check (LDPC) codes maintain a constant encoding ratio [6], which is precisely the regime where RAQM offers the greatest advantage.
This scenario can be understood practically. In an experimental setup with a fixed number of control lines—and thus a fixed number of physical qubits—using a code with constant encoding ratio yields only a fixed number of logical qubits. Changing the code distance does not alter this limit: both the total number of physical and logical qubits remain fixed by the available control lines. By contrast, incorporating memory allows the number of logical qubits to scale despite fixed control resources. Attaching a size-$N$ memory to each physical qubit enables roughly an $N$-fold increase in logical qubits, provided memory read/write errors remain below threshold.
L.2 Case 2: Surface Code
Let $t_{\text{log}}$ and $t_{\text{swap}}$ denote the error correction cycle time and RAQM read/write time. For simplicity, we further assume that the only error channel is decoherence, and all gates (including measurements) are implemented perfectly. We assume that the processor qubit and RAQM memory coherence time are separately $T_{q}$ and $T_{m}$. Without using RAQM, the surface code distance $d_{1}$ should satisfy
$\left(\frac{p_{q}}{p^{\text{th}}_{q}}\right)^{\frac{d_{1}+1}{2}}\leq\varepsilon_{\text{FT}}$ (37)
Here the physical error per correction cycle $p_{q}$ is approximated as $t_{\text{log}}/T_{q}$. $p^{\text{th}}_{q}$ is the error correction threshold for qubit errors. Since we only consider the decoherence error, the two-qubit gate error has the highest weight in the threshold [1], and we approximate $p^{\text{th}}_{q}$ to be $1\%$. To control $k_{0}$ logical qubits with target fault tolerance, the minimum number of control lines per logical qubit is around
$2k_{0}d_{1}^{2}C_{q}.$ (38)
When an RAQM of size $N$ is introduced, the additional physical error rate $p_{m}$ per correction cycle due to the memory is:
$p_{m}\approx\frac{(N-1)(2t_{\text{swap}}+t_{\text{log}})}{T_{m}}+2t_{\text{swap}}\times\frac{1}{2}\left(\frac{1}{T_{m}}+\frac{1}{T_{q}}\right).$ (39)
where the first term represents the accumulated memory idle error, and the latter two terms represent the gate error associated with memory read and write. For the gate error, we assume that the state spends equal time in the qubit and the memory mode, and thus the error includes contributions from both. Here, $p_{m}$ errors are analogous to qubit idle errors [1], which carry a different weight than two-qubit gate errors and therefore correspond to a different error-correction threshold. Here, we take $p^{\text{th}}_{m}=16\%$ as the threshold for memory errors. We note that this value is decoder-dependent and idealized, but the precise number does not affect the discussion below. The minimum surface code distance $d_{2}$ should satisfy:
$\left(\frac{p_{q}}{p^{\text{th}}_{q}}+\frac{p_{m}}{p^{\text{th}}_{m}}\right)^{\frac{d_{2}+1}{2}}\leq\varepsilon_{\text{FT}}$ (40)
and the minimum number of control lines per logical qubit is
$2k_{0}d_{2}^{2}(C_{q}+1)/N.$ (41)
Figure 16: Approximated memory size (left) that maximizes reduction in control lines, and thus yields the highest corresponding memory-less to memory-full control line ratio (right), demonstrating RAQM advantage with the surface code. The approximation is based on the parameters $\{t_{\text{log}},t_{\text{swap}},{p^{\text{th}}_{q}},{p^{\text{th}}_{m}},C_{q}\}=\{200\,\text{ns},500\,\text{ns},1\%,16\%,3\}$ which are slightly pessimistic towards RAQM. Note that the current state-of-the-art parameters are $t_{\text{log}}\approx 1\mathrm{\mu s}$ [1] and $t_{\text{swap}}\approx 100\mathrm{ns}$ [39]. When the RAQM size becomes sufficiently large, additional couplers may be required.
When comparing the control lines in memory-less (Eq. 38) and memory-full (Eq. 41) cases, there exists both a lower and an upper bound to RAQM size for achieving an advantage with the surface code. Intuitively, if the memory contains too few modes, the higher error rate of the swap gate outweighs the benefit of increased storage coherence. Conversely, if the memory is too large, each mode spends considerable time idling while others are being processed, and correcting the resulting accumulated idling errors may require more lines than in the memory-less case.
For given coherence times of the qubit and the memory, we plot both the optimal RAQM size $N$ and control lines ratio (memory-less/memory-full) in Supplemental Figure 16. As shown in the figure, achieving a RAQM advantage requires the memory’s coherence time to be approximately $1$ to $2$ orders of magnitude longer than that of the qubit processor. Using state-of-the-art coherence times $T_{q}=0.1\mathrm{ms}$ and $T_{m}=10\mathrm{ms}$, we can use around $200$ modes in the RAQM to control the same amount of logical qubits with the same logical error rate, while needing $50$ times fewer control lines. As the RAQM size increases, additional couplers may be required, so that the ultimate scaling may vary slightly depending on the specific implementation.
Additionally, unlike classical RAM, where computations slow down due to the extra time required to transfer data between memory and processor, RAQM does not have a similar disadvantage in logical computation speed when heavy error correction code is needed. When performing logical entangling gates in surface code, lattice surgery takes $\text{O}(d_{1})$ code cycles [24]. With transversal entangling gates enabled by RAQM, the code cycle cost scales as $\text{O}(t_{\text{swap}}/t_{\text{log}})$, which is not necessarily greater than that in the memory-less case.
Appendix M Device Fabrication and Measurement Setup